# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ II-C Verifiable machine learning
As machine learning continues to address increasingly complex problems, the size and complexity of the models employed have grown significantly. Larger neural networks require more parameters, which in turn leads to higher costs. These costs stem from the need for extensive datasets and substantial computational resources. As a result, high-cost machine learning computations are often accessible only to large institutions. Within the framework of MLaaS, these institutions, acting as ML service providers, perform ML tasks such as model training or inference and subsequently share the results with their clients.
In this context, it becomes crucial for ML service providers to prove that the shared results were derived from genuine computations rather than being fabricated. This necessity has given rise to the concept of verifiable machine learning. Verifiable machine learning encompasses three primary categories: Verifiable Training, Verifiable Testing, and Verifiable Inference, as outlined below:
  • Verifiable Training ensures the quality of data, the consistency of training algorithms, and the integrity of model parameters throughout the training process. In practice, many individuals and small companies lack the necessary infrastructure to train the machine learning models they require. To address this gap, ML service providers on MLaaS platforms offer model training services, allowing these individuals and small businesses—who act as their ML clients—to outsource their training tasks to the providers on these platforms. In such a setup, the ML service provider undertakes the training process based on the client’s specified configuration details, including accuracy thresholds, the number of epochs, and the network architecture. Once the training is complete, the ML service provider provides the company with the trained model in exchange for a fee. However, this arrangement introduces the need for the individuals and small companies to verify that the ML service provider performed the training task faithfully—strictly adhering to the predefined requirements—and that the returned model is genuinely the result of the stated training process.
  • Verifiable Testing involves proving the true performance of a model, ensuring that its claimed performance accurately reflects its generalization ability rather than being limited to its training data. For instance, in verifiable testing on an MLaaS platform, an ML client uploads the target machine learning model along with some test data and specifies the evaluation metrics to be used, such as accuracy or F1 score. The ML service provider then performs the testing as required and generates a detailed test report. To guarantee the authenticity of this process, some encryption techniques and third-party verification tools must be employed to ensure the following: the test data remains unaltered throughout the testing process; the testing adheres strictly to the predefined configuration; and the final testing results accurately reflect the model’s true performance.
  • Verifiable Inference ensures that the claimed inference results are indeed the outputs generated by the specified machine learning model using the provided input data and following the predetermined inference process. For example, an ML client can upload the input data and the designated model to an ML service provider who performs the inference task while preserving the confidentiality of both the data and the model. To maintain authenticity and data privacy, cryptographic techniques like ZKP can be used. These techniques verify the correctness of the inference process and the integrity of the data without revealing sensitive details. The inference results returned by the ML service provider can then be verified to ensure they have not been tampered with and accurately reflect the model’s capabilities. This approach allows ML clients to safely and reliably utilize external ML inference services.
@@
TABLE II: Comparison of Security Techniques in Verifiable Machine Learning
Security Techniques Advantages in Verifiable Machine Learning Limitations Compared to ZKP
DP [22] Protects training data from inference attacks by introducing statistical noise The added noise reduces dataset utility, impacting model accuracy and usability
HE [23] Enables computations on encrypted data, ensuring data privacy High computational and storage overheads limit scalability and efficiency
FL [24] Allows collaborative training without data sharing, protecting privacy and reducing overall computational costs Vulnerable to malicious participants providing false data, which can compromise the model
TEE [25] Isolates sensitive computations in a secure environment with integrity and confidentiality guarantees Unsuitable for tasks requiring extensive computational resources
SMC [26] Achieves high computational accuracy without leaking participants’ private data; supports general-purpose computations Computational delay does not scale linearly with the number of participants, affecting efficiency
II-D Comparison of Security Techniques in Verifiable Machine Learning
Several security techniques, such as Differential Privacy (DP), Homomorphic Encryption (HE), Federated Learning (FL), Trusted Execution Environments (TEE), and Secure Multiparty Computation (SMC), can provide verifiable computation and privacy protection in machine learning to a certain degree. However, when these security techniques are applied to verifiable machine learning, they exhibit certain limitations compared to ZKP. The comparative analysis is presented in Table II and detailed as follows.
DP is a highly regarded and rigorous security technique for privacy protection. Initially proposed by Dwork in 2006 [22], DP operates by adding random noise to query results, ensuring that the presence or absence of any individual in the dataset does not significantly alter the outcome. This approach safeguards individual data privacy. When applied to ML models, DP can protect training data against model inversion attacks. Consequently, numerous studies have explored integrating DP into ML models [27]. However, compared to ZKP, the introduction of random noise can compromise dataset accuracy, potentially affecting both model utility and precision.
HE [23] is an encryption technique that enables computations to be performed directly on encrypted data (ciphertexts). The results remain encrypted, and once decrypted, they match the outcomes of the same computations performed on plaintexts. This capability is often referred to as ”computable yet invisible” data. HE is categorized into Fully Homomorphic Encryption (FHE) and Partially Homomorphic Encryption (PHE). FHE supports arbitrary computations on ciphertexts, whereas PHE allows only specific operations, such as addition, multiplication, or a limited combination of both. While HE ensures data privacy while maintaining data computability, it faces significant performance challenges, including high computational and storage overhead, especially when compared to ZKP.
FL is a distributed machine learning framework introduced by Google in 2016 [24, 28]. It enables collaborative model training while preserving data privacy and ensuring regulatory compliance, thereby enhancing AI model performance. FL addresses the limitations of single-feature data during the training phase and safeguards private data against leakage. Additionally, its distributed architecture reduces overall computational costs. However, FL has a significant vulnerability: it cannot prevent participants from submitting false or malicious data, which can irreparably compromise the final trained model.
TEE [25] is a secure, independent processing environment with computation and storage capabilities, designed to provide robust security and integrity protection. It utilizes isolated memory to store private data and perform computations, ensuring that only authorized interfaces can access the data. Hardware-based TEE technology is highly efficient and capable of supporting multi-level, complex algorithm logic implementations. However, its reliance on underlying hardware architecture makes it less suitable for tasks requiring high network bandwidth or significant computational resources.
SMC [26] is a cryptographic technique that enables multiple parties to collaboratively achieve computational goals while ensuring that private data remains confidential, except for the computed results and any information inferable from them. SMC offers high computational accuracy and supports programmable, general-purpose computations. Nevertheless, as the number of participants increases, it becomes increasingly difficult to design computational schemes that ensure computation latency scales linearly, posing challenges for large-scale applications.
III Research of ZKML
In this section, we begin by introducing verifiable machine learning in the context of zero-knowledge proofs, referred to as ZKML. Next, we categorize and analyze the existing research efforts in ZKML. Finally, we discuss the key challenges faced in the development of ZKML, as well as the potential solutions proposed to address these challenges.
III-A Introduction of ZKML
In ZKML, there are two key participants: the ML service provider, acting as the prover in the ZKP system ($P$), and the ML client, serving as the verifier in the ZKP system ($V$). Due to $V$’s limited computational resources or lack of access to the necessary data, it delegates machine learning tasks—including training and inference—to $P$. As a result, the majority of computations in the machine learning workflow are performed by $P$. Meanwhile, $V$ focuses solely on verifying the correctness of the results and the processes used to produce them, ensuring both accuracy and integrity.
Depending on the specific tasks and privacy protection requirements, the prover $P$ and the verifier $V$ may hold different datasets, models, and objectives. Since both parties operate in a trustless environment, $P$ might act dishonestly—for example, by using incorrect data during training or fabricating seemingly valid results—which could severely compromise the learning task’s outcome. To address these issues, ZKML is introduced as a means to ensure transparency and trustworthiness in the computational process.
The typical workflow of ZKML is illustrated in Fig. 2. First, the prover $P$ and the verifier $V$ agree on the machine learning task $T$ (i.e., to specify the used model and performance metric, etc.) and input data $D$. The verifier $V$ generates and sends a commitment $c$ to the prover $P$ to confirm the integrity of the committed input data $D$. After verifying the commitment, the prover $P$ generates the proving and verification keys. Subsequently, the prover $P$ executes the machine learning task to obtain the result $r$ while simultaneously using a proving mechanism to generate a proof $\pi$. Finally, the verifier $V$ evaluates the result $r$ and the proof $\pi$ to determine whether the prover $P$ has correctly executed the machine learning task $T$. If the verification of $\pi$ is successful, the verifier $V$ concludes that the prover $P$ has honestly performed the task $T$ based on the input data $D$, producing the correct result $r$. Otherwise, the verifier $V$ rejects the result.
Figure 2: The typical workflow of ZKML.
@@ III-A Introduction of ZKML
In a ZKP system, the statement to be proven (e.g., verifying that a computation was performed correctly or that a secret satisfies certain conditions) is encoded as a computation. This computation is typically represented as an arithmetic circuit, which processes public inputs and private witnesses through a sequence of arithmetic operations to generate an output. In this paper, we exploits conceptual diagrams to depict the designs of the arithmetic circuits for the three categories of ZKML. These diagrams highlight three primary components: circuit input, circuit logic, and circuit output. The configuration of these components—circuit input, circuit logic, and circuit output—varies across the three categories of ZKP-based verifiable machine learning, as described below.
III-A1 ZKP based Verifiable Training
@@
  • Circuit Input: The inputs to the circuit in ZKP-based verifiable training include both private witness and public input. The private witness encompasses the training data, the labels of the training data. The public input consists of the hash of the training data, the hash of the training data labels, the threshold of the model training loss and the proving key necessary to generate zero-knowledge proofs. [^2] The hashes used in the inputs serve as commitments to the corresponding data. Alternatively, other cryptographic commitment schemes compatible with ZKP can also be utilized.
  • Circuit Logic: The circuit executes several essential computational tasks to enable verifiable training. First, it computes the hashes of the training data and their corresponding labels that given in the private witness, then compares these computed hashes with those provided in the public input. If the hashes match, the computation proceeds; otherwise, it terminates. Next, the circuit generates model inference outputs using the training data as inputs to the model. It then calculates the training loss by comparing the inferred outputs with the true labels of the training data. This training loss is evaluated against a predetermined threshold. If the loss exceeds the threshold, the circuit computes gradients and updates the model parameters accordingly; otherwise, the training process halts. Throughout this process, all computations are integrated with the proving key to produce a zero-knowledge proof, ensuring the integrity and confidentiality of the entire computation process.
  • Circuit Output: The output includes the hash of the updated model parameters after training and the generated zero-knowledge proof.
This circuit design enables the verification of each step in the training process according to the defined algorithms and procedures, without revealing sensitive information such as the training data and the trained model parameters. Consequently, it ensures the transparency and trustworthiness of the training process while preserving data and model privacy.
Figure 3: Conceptual diagram of the arithmetic circuits used in ZKP-based verifiable training.
III-A2 ZKP based Verifiable Testing
@@
  • Circuit Input: Within the input to the circuit, the private witness includes the model parameters; the public input consists of the the testing data and their corresponding labels, the hash of the model parameters, the threshold of the model testing performance and the proving key necessary to generate zero-knowledge proofs.
  • Circuit Logic: The circuit performs several key computational tasks for verifiable testing. First, it computes the hash of the model parameters, then compares the computed hash with the hash provided in the public input. If the hashes match, the computation proceeds; otherwise, it terminates. After that, it computes the model inference outputs using the testing data as the inputs to the model with the model parameters. Next, it calculates the model testing performance by comparing the inferred outputs with the true labels of the testing data. This computed model testing performance is then evaluated against a predetermined performance threshold. If the performance exceeds the threshold, a zero-knowledge proof for this result is generated using the proving key.
  • Circuit Output: The output includes the testing performance of the model and the generated zero-knowledge proof.
This circuit design of ZKP-based verifiable testing ensures that the model testing performance is within a certain range, meanwhile the privacy of the model parameters is protected.
Figure 4: Conceptual diagram of the arithmetic circuits used in ZKP-based verifiable testing.
@@ III-A3 ZKP-based Verifiable Inference
  • Circuit Input: The inputs to the circuit in ZKP-based verifiable inference include both private witness and public input. The private witness comprises the model parameters. The public input includes the hash of the model parameters, the data used for model inference, and the proof key required for generating the zero-knowledge proof.
  • Circuit Logic: The circuit performs several critical computational tasks for verifiable inference. It first computes the hash of the model parameters, then compares the computed hash with the hash provided in the public input. If the hashes match, the computation proceeds; otherwise, it terminates. Then, it computes the inferred outputs by applying the model parameters to the input data. Simultaneously, the proof key is utilized to generate a zero-knowledge proof of the correctness of the inference computation.
  • Circuit Output: The output consists of the inference results for the input data and the generated zero-knowledge proof.
This circuit design ensures that the inference process adheres to the defined algorithms and procedures, allowing for verification of correctness without exposing sensitive information (i.e., the model parameters). Consequently, it upholds the privacy of the model parameters while maintaining the integrity and trustworthiness of the inference results.
Figure 5: Conceptual diagram of the arithmetic circuits used in ZKP-based verifiable inference.
@@
III-B Discussion of Existing ZKML Studies
Since its inception in 2017, research in ZKML has progressed rapidly, with a significant body of work emerging by 2024. Fig. 6 presents a timeline of representative studies, highlighting the advancements and evolution in this field. These studies can be categorized into the three primary categories of verifiable machine learning (i.e., verifiable training, verifiable testing, and verifiable inference) based on the stages of the machine learning process and their verification objectives. The detailed classifications and corresponding research outcomes are summarized in Table III. In the following, we examine the works of each category in depth, analyzing the design principles, implementation methodologies, and application scenarios of various approaches. This discussion aims to shed light on how these methods enhance the security and trustworthiness of machine learning tasks.
III-B1 ZKP based Verifiable Training

### v2

@@ II-C Verifiable Machine Learning
As ML models become larger and more expensive, outsourcing training, testing, and inference to external platforms has become increasingly common. In such settings, a client typically lacks the resources required to repeat the entire computation locally and therefore needs a mechanism to verify the integrity of the outsourced result. Verifiable machine learning studies how a service provider can furnish convincing evidence that an ML task has been executed according to a declared dataset, model, and evaluation procedure.
In the ZKML setting, the service provider typically acts as the prover $\mathcal{P}$, and the client acts as the verifier $\mathcal{V}$. Depending on the stage of the ML pipeline being certified, verifiable machine learning can be divided into three primary categories: Verifiable Training, Verifiable Testing, and Verifiable Inference.
  • Verifiable Training certifies that the updated model parameters $\theta^{\prime}$ are obtained by running the declared training procedure on the committed training data under the agreed hyperparameters and architecture. This category is relevant when a client outsources model training but still needs evidence that the delivered model is the genuine result of the specified optimization process rather than a fabricated or improperly trained artifact.
  • Verifiable Testing certifies that a reported evaluation metric, such as accuracy or F1 score, is computed from the declared test dataset and the committed model under a specified evaluation protocol. Its goal is to prevent misleading claims about model quality, for example when a provider reports inflated performance that does not correspond to the agreed test setting.
  • Verifiable Inference certifies that the returned prediction $\hat{y}$ is exactly the output produced by the specified model on the declared input $x$. This category is especially important in MLaaS scenarios where the client wants assurance that the inference result is genuine, while the provider wants to keep the model parameters confidential.
@@
TABLE III: Comparison of representative security techniques for verifiable machine learning.
Technique Security Basis Primary Security Property Runtime Overhead Verification Capability
DP Statistical privacy model Record-level privacy via randomized noise Low None
HE Cryptographic hardness assumptions Confidential computation on encrypted data Very high None
FL Trust in protocol participants / coordinator Data locality and collaborative training Low to moderate None
TEE Trust in hardware vendor and platform Hardware-backed execution isolation Low Limited (attestation)
SMC Cryptographic assumptions Multi-party input privacy High Limited
ZKP Cryptographic assumptions Privacy plus verifiable computation Moderate to high Strong
II-D Comparison of Security Techniques for Verifiable Machine Learning
Several security techniques, including Differential Privacy (DP), Homomorphic Encryption (HE), Federated Learning (FL), Trusted Execution Environments (TEE), and Secure Multiparty Computation (SMC), are important for privacy-preserving and trustworthy ML. However, they target different threat models and provide different guarantees. In particular, none of them, when used alone, offers the same combination of privacy preservation and succinct public verifiability as ZKP. Table III summarizes the comparison, and the main distinctions are outlined below.
DP is a rigorous privacy framework that protects individual records by injecting calibrated random noise [19, 20]. In ML, DP is widely used to reduce the leakage risk of training data. However, the added noise may reduce model utility, and DP does not provide any mechanism for proving that the claimed privacy budget or computational procedure was actually enforced.
HE enables direct computation on encrypted data [21]. This makes HE highly attractive for confidential inference and secure outsourced computation. Its main limitation is efficiency: ciphertext expansion, noise management, and expensive homomorphic operations often introduce substantial computational and storage overhead. Moreover, HE alone does not let a client verify that the service provider executed the encrypted computation correctly.
FL keeps raw data local and supports collaborative model training across multiple participants [22, 23]. This data-locality property can improve privacy and regulatory compliance. Nevertheless, FL does not natively prevent malicious participants from submitting poisoned or low-quality updates, and it offers no inherent cryptographic proof that the aggregation or local training steps were executed correctly.
TEE provides hardware-backed isolation for storing data and executing code securely [24]. Compared with purely cryptographic techniques, TEE is usually efficient and can support complex programs with low runtime overhead. Its limitations stem from its trust model: users must trust the hardware vendor and platform implementation, and the resulting security can be weakened by side-channel attacks, limited enclave memory, and implementation vulnerabilities. Although remote attestation offers some assurance about the execution environment, it is not equivalent to a succinct public proof of each computation step.
SMC enables multiple parties to jointly compute a function while keeping their private inputs hidden [25]. It offers strong privacy guarantees and is well suited to collaborative settings. However, SMC protocols often incur high communication overhead, multiple interaction rounds, and significant implementation complexity, which makes large-scale ML deployment challenging. In addition, the resulting guarantees are usually protocol-specific rather than succinctly verifiable by an arbitrary third party.
Overall, these techniques should be viewed as complementary rather than mutually exclusive. ZKP is distinctive because it directly addresses computational verifiability while still protecting private witnesses. This makes it a natural foundation for verifiable ML, and it also explains why recent systems increasingly combine ZKP with HE, FL, DP, or TEE to obtain stronger end-to-end guarantees. Section V revisits these hybrid designs in more detail.
Figure 2: Typical workflow of a ZKML protocol. The verifier $\mathcal{V}$ and prover $\mathcal{P}$ first establish a shared task agreement specifying the task $T$, the circuit $C_{T}$, the public input $x$, and the witness $w$. The verifier commits to the relevant information through a digest $c=h(\cdot)$, while the setup procedure generates $(\mathsf{PK},\mathsf{VK})$. The prover then checks the commitment, executes the delegated ML task on $(x,w)$, obtains the result $r$, and generates a proof $\pi$. Finally, the verifier runs $\mathsf{VERIFY}(\mathsf{VK},x,\pi)$ and outputs an accept/reject decision.
III ZKML Research Landscape
This section reviews the research landscape of ZKML. We first outline the general workflow and design logic of ZKML systems, then survey representative studies across verifiable training, testing, and inference, and finally summarize the main implementation challenges together with the technical improvements proposed to address them.
III-A ZKML Overview
In ZKML, there are two key participants: the ML service provider acts as the prover in the ZKP system, denoted by $\mathcal{P}$, and the ML client acts as the verifier, denoted by $\mathcal{V}$. Because $\mathcal{V}$ may have limited computational resources or limited access to the required data and models, it delegates computationally intensive ML tasks, such as training, testing, or inference, to $\mathcal{P}$. Accordingly, most of the ML computation is carried out by $\mathcal{P}$, while $\mathcal{V}$ focuses on checking the correctness and integrity of the returned result.
Depending on the task setting and privacy requirements, $\mathcal{P}$ and $\mathcal{V}$ may possess different datasets, models, and verification objectives. Since both parties operate in a trustless environment, a dishonest prover may deviate from the agreed workflow, for example by using incorrect data, skipping part of the computation, or fabricating a seemingly valid result. ZKML addresses this problem by attaching cryptographic verifiability to the ML computation.
The typical workflow of ZKML is illustrated in Fig. 2. First, the prover $\mathcal{P}$ and the verifier $\mathcal{V}$ agree on a task instance $T$, the corresponding arithmetic circuit $C_{T}$, and the public information to be exposed during verification. The verifier $\mathcal{V}$ commits to the relevant data or model state through a digest $c=h(\cdot)$ and sends this commitment to $\mathcal{P}$. After the setup phase generates $(\mathsf{PK},\mathsf{VK})\leftarrow\mathsf{KEYGEN}(1^{\lambda},C_{T})$, the prover $\mathcal{P}$ executes the ML task, obtains the result $r$, and produces a proof $\pi\leftarrow\mathsf{PROVE}(\mathsf{PK},x,w)$, where $x$ and $w$ denote the public input and private witness, respectively. Finally, the verifier $\mathcal{V}$ checks the claimed result together with $\pi$ by running $\mathsf{VERIFY}(\mathsf{VK},x,\pi)$. If verification succeeds, $\mathcal{V}$ accepts that $\mathcal{P}$ has correctly executed $T$; otherwise, it rejects the result.
@@ III-A ZKML Overview
In a ZKP system, the statement to be proven (e.g., verifying that a computation was performed correctly or that a secret satisfies certain conditions) is encoded as a computation. This computation is typically represented as an arithmetic circuit, which processes public inputs and private witnesses through a sequence of arithmetic operations to generate an output. In the following discussion, we use $h(\cdot)$ to denote a hash-based commitment, $\mathcal{D}_{\mathrm{train}}=(X_{\mathrm{train}},Y_{\mathrm{train}})$ and $\mathcal{D}_{\mathrm{test}}=(X_{\mathrm{test}},Y_{\mathrm{test}})$ to denote the training and testing datasets, $\theta$ to denote model parameters, $\mathsf{PK}$ and $\mathsf{VK}$ to denote the proving and verification keys, and $\pi$ to denote the generated proof. The conceptual diagrams in this paper highlight three primary components: circuit input, circuit logic, and circuit output. Their configurations vary across the three categories of ZKP-based verifiable machine learning, as described below.
III-A1 ZKP-based Verifiable Training
@@
  • Circuit Input: The circuit takes both a private witness and a public input. The witness contains the training features and labels, denoted by $(X_{\mathrm{train}},Y_{\mathrm{train}})$. The public input contains the commitments $h(X_{\mathrm{train}})$ and $h(Y_{\mathrm{train}})$, the target loss threshold $\tau_{\mathrm{loss}}$, and the proving key $\mathsf{PK}$. [^2] The hashes used in the inputs serve as commitments to the corresponding data. Alternatively, other cryptographic commitment schemes compatible with ZKP can also be utilized.
  • Circuit Logic: The circuit first recomputes $h(X_{\mathrm{train}})$ and $h(Y_{\mathrm{train}})$ from the witness and checks them against the public commitments. If the commitments are consistent, it evaluates the model on $X_{\mathrm{train}}$, computes the training loss against $Y_{\mathrm{train}}$, and compares the resulting loss with $\tau_{\mathrm{loss}}$. If the loss remains above the threshold, the circuit computes gradients and updates the model parameters; otherwise, the training process halts. The entire computation is coupled with $\mathsf{PK}$ to generate a zero-knowledge proof $\pi$, thereby ensuring integrity without revealing the private training data.
  • Circuit Output: The output includes the commitment $h(\theta^{\prime})$ of the updated model parameters $\theta^{\prime}$ and the generated proof $\pi$.
This circuit design enables $\mathcal{V}$ to verify that the updated model commitment $h(\theta^{\prime})$ is obtained from the declared training workflow and the committed training dataset, without revealing the private training data or model parameters. Consequently, it ensures the transparency and trustworthiness of the training process while preserving data and model privacy.
Figure 3: Conceptual arithmetic-circuit design for ZKP-based verifiable training. The figure separates the computation into circuit input, circuit logic, and circuit output. The public inputs include the hashes of the training data and labels, the loss threshold, and the proving key, whereas the private witness contains the underlying training data and labels. Inside the circuit, the hashes are recomputed and checked for consistency, after which the generated training output is compared with the ground-truth labels to evaluate whether the declared loss condition is satisfied. The circuit output shown in the figure consists of the hash of the updated model together with a true/false validity outcome.
III-A2 ZKP-based Verifiable Testing
@@
  • Circuit Input: The private witness contains the model parameters $\theta$. The public input contains the testing dataset $\mathcal{D}_{\mathrm{test}}=(X_{\mathrm{test}},Y_{\mathrm{test}})$, the parameter commitment $h(\theta)$, the performance threshold $\tau_{\mathrm{perf}}$, and the proving key $\mathsf{PK}$.
  • Circuit Logic: The circuit first recomputes $h(\theta)$ from the witness and compares it with the public commitment. If the commitments match, it evaluates the model on $X_{\mathrm{test}}$, computes the testing metric $m_{\mathrm{test}}$ against $Y_{\mathrm{test}}$, and compares the resulting score with $\tau_{\mathrm{perf}}$. If the performance requirement is satisfied, the circuit uses $\mathsf{PK}$ to generate the proof $\pi$.
  • Circuit Output: The output includes the testing metric $m_{\mathrm{test}}$ and the generated proof $\pi$.
This circuit design enables $\mathcal{V}$ to verify that the reported testing metric $m_{\mathrm{test}}$ is computed from the committed model $\theta$ and the declared testing dataset $\mathcal{D}_{\mathrm{test}}$, while keeping the model parameters hidden.
@@ III-A3 ZKP-based Verifiable Inference
  • Circuit Input: The inputs to the circuit in ZKP-based verifiable inference include both a private witness and a public input. The witness contains the model parameters $\theta$. The public input includes the parameter commitment $h(\theta)$, the inference input $x$, and the proving key $\mathsf{PK}$.
  • Circuit Logic: The circuit first recomputes $h(\theta)$ from the witness and compares it with the public commitment. If the commitments match, it evaluates the model on the input $x$ to obtain the prediction $\hat{y}=f_{\theta}(x)$. At the same time, $\mathsf{PK}$ is used to generate a zero-knowledge proof $\pi$ for the correctness of the inference computation.
  • Circuit Output: The output consists of the prediction $\hat{y}$ and the generated proof $\pi$.
This circuit design enables $\mathcal{V}$ to verify that the prediction $\hat{y}$ is generated by the committed model $\theta$ on the declared input $x$, without exposing the model parameters. Consequently, it preserves model privacy while maintaining the integrity and trustworthiness of the inference result.
@@
Figure 4: Conceptual arithmetic-circuit design for ZKP-based verifiable testing. The figure shows a three-part view consisting of circuit input, circuit logic, and circuit output. The public input contains the hash of the model parameters, the testing data and corresponding labels, the performance threshold, and the proving key, while the private witness contains the model parameters themselves. The circuit recomputes the parameter hash, checks consistency with the public commitment, evaluates the model on the declared test set, and compares the resulting testing performance against the stated threshold. The circuit output shown here is the testing performance of the model together with a true/false validity outcome.
Figure 5: Conceptual arithmetic-circuit design for ZKP-based verifiable inference. The figure shows that the public input contains the hash of the model parameters, the data for model inference, and the proving key, whereas the private witness contains the model parameters themselves. Inside the circuit, the parameter hash is recomputed and checked against the public commitment before the model is applied to the input data to compute the inference result. The circuit output shown in the figure is the inference result together with a true/false validity outcome, allowing correctness to be certified without revealing the private model parameters.
III-B Representative ZKML Studies
Since the emergence of SafetyNets in 2017, ZKML has evolved from early proofs of concept into a diverse research landscape spanning training, testing, and inference verification. Fig. 6 outlines the timeline of representative studies. For clarity, we organize these works by verification objective, namely verifiable training, verifiable testing, and verifiable inference. Their overall classification is summarized in Table VII. To facilitate direct efficiency comparisons, we further report proving time, verification time, and proof size in three dedicated quantitative tables: Table IV for verifiable training, Table V for verifiable testing, and Table VI for verifiable inference. The discussion below focuses on the core design ideas, implementation strategies, and application settings of representative systems, thereby clarifying how ZKML has progressed toward more trustworthy and practical machine learning services.
Figure 6: Timeline of representative ZKML studies. The figure highlights the evolution of the field from early proof-of-concept systems for neural-network inference and decision-tree verification to more recent work on verifiable training, commit-and-prove optimization, large-model inference, and LLM-oriented systems. It also illustrates how the research focus gradually shifted from small and highly specialized models toward broader system support, higher efficiency, and stronger deployment realism.
TABLE IV: Quantitative comparison of representative ZKP-based verifiable training systems.
Work Model(s) Hardware / Platform Proving Time Verification Time Proof Size
VeriML [26] Six models (e.g., NN) Intel Core i5-4460S <20<20s (e.g., 12s for NN, batch size 32) <2<2s (e.g., 0.99s for NN) N/A
zkMLaaS [27] LogiR 24 Intel Xeon E5-262, 8 V100 GPUs, 128GB RAM 2.2s 5ms 24.1MB
ZKDL [28] DNN 12 CPU cores, Tesla A100 GPU, 128 GB RAM 0.86s (1024 samples, batch size 64) 0.19s 0.033KB
ZKPoT [29] LogiR N2 GCP, 512GB RAM 4208s (total) 26.5s (total) <350<350MB
Kaizen [30] VGG-11 (10M parameters) 8x Xeon Platinum 8370, 512GB RAM <882<882s (per sample, batch size 16) <0.130<0.130s <1627<1627KB
VeriLoRA [31] LLM (<<13B, e.g., LLaMA-2) AMD EPYC 9654, Nvidia A100 (80GB), 500GB RAM <600<600s (per sample, batch size 1) <4<4s N/A
TABLE V: Quantitative comparison of representative ZKP-based verifiable testing systems.
Work Model(s) Hardware / Platform Proving Time Verification Time Proof Size
ZKDT [32] DT (1029 nodes, 23 levels, 5000 samples) Intel Xeon Platinum 8124M 250s (accuracy testing) 15.6s 287KB
pvCNN [33] CNN (e.g., LeNet-5) Quad-core i5-7500, 48GB RAM 1448.83s (ReLU matrix) 54s (MNIST) 351MB
ZEN [34] NN (ShallowNet, LeNet) 32-core Xeon 8280M, 256GB RAM 147s to 4710s 0.023s to 0.47s 192 bytes
zkCNN [35] CNN AMD EPYC 7R32, 128GB RAM 88.3s 59.3ms 341KB
Artemis [36] ML (GPT-2) 128 vCPUs, 1TB RAM 200-240 min (GPT-2) 14ms (GPT-2) 15KB (GPT-2)
TABLE VI: Quantitative comparison of representative ZKP-based verifiable inference systems.
Work Model(s) Hardware / Platform Proving Time Verification Time Proof Size
Drynx [37] Regression models Intel Xeon E5-2680 v3 CPUs, 256GB RAM <22<22s (LogR, 600k records) <22<22s 56kB
vCNN [38] CNNs (VGG16) Quad-core Intel CPU i5 8 hours (VGG16) 19.4s (VGG16) N/A
Mystique [39] NN (ResNet-101, 42.5M parameters) m5.2xlarge, 32GB RAM 262s N/A 0.99GB
Singh [40] DT 8x Xeon, 32GB RAM 170-200s 0.4s N/A
ZKML [41] ML (GPT-2) 128 vCPUs, 1TB RAM 3652s (GPT-2) 18.7s (GPT-2) 28KB (GPT-2)
zkLLM [42] LLM (<<13B, e.g., LLaMA-2) Nvidia A100 (40GB), 124GB RAM <15<15 min <3<3s <10<10KB
Lu [43] NN (GPT-2, 117M parameters) Ryzen 3700X 8-core, 32GB RAM 287.1s (GPT-2) 63s (GPT-2) 4.5GB (GPT-2)
Hao [44] LLMs AWS c5.9xlarge 2.1s (ReLU); 87s (Softmax) N/A 30MB (ReLU); 816MB (Softmax)
III-B1 ZKP-based Verifiable Training